OpenStudy (anonymous):

Together the Yankees and Red Sox have won 30 World Series in the last 100 years. The Yankees have won 9 times as many as the Red Sox. How many World Series have the Yankees won? Let the Yankees be represented by x and the Red Sox by y. 27World Series 21World Series 3 World Series 15World Series

OpenStudy (anonymous):

hmm?

OpenStudy (zzr0ck3r):

Let \(y\) be the amount of games the Yankees one. Let \(r\) be the amount of games the Red Sox one. then they say the sum of these games is \(30\) so \(x+y=30\) the Yankees won 9 times as many games so if \(x=1 y = 9\) if \(x = 2 y = 18\) in other words \(y=9x\) we plug \(\uparrow\) this into the other equation for \(y\) \(x+y=30\) becomes \(x+9x=30\) can you solve for \(x\)?

OpenStudy (zzr0ck3r):

change \(r\) to \(x\) is the second line